OpenStudy (anonymous):

write an equation for the line that is parallel to the given line & passes through the given point. y=x+8;(2,16)

OpenStudy (afifarahman):

\[x-y+10=0\]

OpenStudy (anonymous):

It's multiple choice A) y=x+14 By=x C) y=x+6 D) y=x-6

OpenStudy (anonymous):

I got it y=x+14